Manuel invested his savings in two investment funds. The amount he invested in Fund A was
$
2000
less than the amount he invested in Fund B. Fund A returned a
5
%
profit and Fund B returned a
7
%
profit. How much did he invest in Fund A, if the total profit from the two funds together was
$
1100
?

Answer by mananth(16946)

B -Amount @ 7.00% x
A- Amount @ 5.00% 1 x -2000
Total interest earned = 1100

7.00% x+ 5.00% ( 1 x -2000 )= 1100
7.00% x+ 5.00% * 1 x -2000 = 1100
7.00% x+ 5 x -10000 = 110000
7.00 x+ 5 x = 110000
12.00 x = 120000
x= 10000 Amount @ 7.00%

Amount @ 5.00% = 10000 -2000
= 8000

Interest
CHECK
700 + 400 = 1100